OpenStudy (anonymous):

4(7 - 3x) = 7(4 - 2x)

OpenStudy (lieutenantgeneral):

x = 0

OpenStudy (anonymous):

x=0

OpenStudy (jadeishere):

28 - 12x = 28 - 14x 2x = 0 x = 0